1.A certain bank pays 1% compound interest for their savings account holders. If Nick deposits $10000 in that bank. (refer page 707 in the text book) (40 points) Formula = A= P (1 + r)^t
a. Estimate how much money will be in his account after 25 years using your formula. Assume there will not be any deposits or withdrawals during that 25 year period.

b. Estimate how long it will take to double the total amount (i. e. the total is $20000) in Nick’s account.

c.Estimate the total amount in Nick’s account after one year.

Answer by stanbon(75887)

A certain bank pays 1% compound interest for their savings account holders. If Nick deposits $10000 in that bank. (refer page 707 in the text book) (40 points) Formula = A= P (1 + r)^t
a. Estimate how much money will be in his account after 25 years using your formula. Assume there will not be any deposits or withdrawals during that 25 year period.
A(25) = 10,000(1+0.01)^25
A(25) = 10,000(1.01)^25
A(25) = 10,000*1.2824
A(25) = 12,824.00
=======================
b. Estimate how long it will take to double the total amount (i. e. the total is $20000) in Nick’s account.
20,000 = 10,000(1.01)^t
2 = 1.01^t
t*log(1.01) = log(2)
t = log(2)/log(1.01)
t = 69.66 years
==========================
c.Estimate the total amount in Nick’s account after one year.
A(1) = 10,000(1.01)^1
A(1) = 10,100
=====================
Cheers,
Stan H.